# Symplectic mapping class relations from pencil pairs

### v1

@@ Abstract.
We describe symplectic mapping class relations between products of positive Dehn twists along Lagrangian spheres in Weinstein $4$-manifolds, all of which are affine $\mathbb{C}$ varieties. The relations are obtained by applying classification results for Fano $3$-folds and polarized $K_{3}$ surfaces of small genus to a general methodology – finding pencil pairs.
@@ 1. Introduction
This article describes a general method for finding symplectic mapping class relations between products of positive Dehn twists on (affine algebraic) Weinstein manifolds $F$ of $\dim=2n$ by studying complex projective varieties of $\dim_{\mathbb{C}}=n+1$. The general method consists of finding pencil pairs and is known to experts. It recovers known mapping class relations on punctured Riemann surfaces such as the lantern relation [AS08, J79], described in §3.3 and depicted in Figure 1. In formally defining pencil pairs and working out some basic techniques for finding them (see §4), we hope to inspire specialists in symplectic topology and algebraic geometry to further study the case $\dim F\geq 4$, for which few examples of such symplectic mapping class relations are known. Therefore our exposition will cover some relevant basics from both subject areas. Our search for mapping class relations on high-dimensional Weinstein manifolds is motivated by recent advances in high-dimensional contact topology. See §2.3.
@@
Figure 1. The lantern relation states that a product of Dehn twists along circles shown on the left is isotopic (relative to the boundary of the $4$-holed sphere) to a product of Dehn twists along the circles shown on the right.
In [O21] Oba finds a symplectic mapping class relation on a Weinstein $4$-manifold by applying the classification of del Pezzo surfaces [F80] to find a pencil pair. Our main application applies classification results for polarized $K_{3}$ varieties [M87, SD74] and Fano $3$-folds [I77, I77, MM82] to find new pencil pairs, and so new examples of mapping class relations on Weinstein $4$-manifolds. Here is one such family of examples.
@@
Let $Z\subset\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{2}$ be the $K_{3}$ surface cut out by a smooth complete intersection of $\deg=(1,1,2)$ and $\deg=(1,1,1)$ divisors. Let $(F,\beta_{F})$ be the Weinstein manifold given by the complement of a $\deg=(1,1,1)$ divisor in $Z$, with the symplectic structure determined by a $\deg=(1,1,1)$ line bundle. Then there is a compactly supported symplectomorphism $\tau_{\partial F}\in\pi_{0}\operatorname{Symp}^{c}(F,d\beta_{F})$ which can be factored into a product of $m$ Dehn twists along Lagrangian spheres in $(F,\beta_{F})$, where
$m=18,32,34,36,40.$
We found six such families of mapping class relations in total, each giving rise to an infinite family of mapping class relations by a generally-applicable “cabling” argument described in Lemma 4.2.1. One of these families (see §5.3) recovers the relation of [O18]. Because the number of Dehn twists in our relations are all distinct, they cannot be derived from well-known braid relations [KS02]. See §2.2.
@@ 1.1. Conventions
All varieties in this article are smooth and defined over $\mathbb{C}$. Throughout $\dim$ and $\dim_{\mathbb{C}}$ refer to real and complex dimension. On a closed manifold $X$ we use $\operatorname{PD}_{X}$ for Poincaré duality in either direction $H_{k}(X)\leftrightarrow H^{\dim X-k}(X)$. All of our symplectic manifolds will be complex projective varieties $X,Z,B,\mathcal{M}\subset\mathbb{P}^{n}$ or affine varieties $F,W\subset\mathbb{C}^{n}$, with their symplectic structures inherited from the ambient space.
1.2. Acknowledgements
We thank Eduardo Alves da Silva, Yankı Lekili, and Takahiro Oba for interesting discussions. We’re especially grateful to Marco Golla for discussions around complex surfaces, reviewing an early draft of the article, and for pointing out the $l>1$ examples of §3.3. We thank Cofund MathInGreaterParis and the Fondation Mathématique Jacques Hadamard for supporting us as a member of the Laboratoire de Mathématiques d’Orsay at Université Paris-Saclay. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034255.
@@ 2.1. Symplectic mapping class relations
An exact Lagrangian sphere $L\subset F$, where $\dim F=2n$ is a sphere $\mathbb{S}^{n}$ for which $\beta_{F}|TL=df$ for some $f\in\mathcal{C}^{\infty}(L)$. A framing on such a sphere is an identification of the sphere with the boundary of a disk, up to homotopy. We give a quick review of Dehn twists about such $L$, which appear in nature as monodromies of Milnor fibers of the singularities $\sum_{0}^{n}z_{j}^{2}$ on $\mathbb{C}^{n+1}$ [Arn95]. According to the Weinstein neighborhood theorem, every such $L$ has a neighborhood of the form $(N_{F}(L),\beta_{F})=(\mathbb{D}^{\ast}\mathbb{S}^{n},pdq+df)\implies d\beta_{F}|_{N_{F}(L)}=dp\wedge dq$ where $pdq$ is the canonical $1$-form on a cotangent disk bundle $\mathbb{D}^{\ast}\mathbb{S}^{n}$ of the sphere and $f\in\mathcal{C}^{\infty}(\mathbb{D}^{\ast}\mathbb{S}^{n})$ is some smooth function. In particular, writing $\bullet$ for the intersection pairing on $H_{n}(L)$, $[L]\bullet[L]$ is 00 for $n$ odd. When $n$ is even, $[L]\bullet[L]=-2$ meaning that $[L]$ must be primitive in $H_{n}(F)$. Let’s assume the disk bundle is defined using the standard round metric of radius $1$. Then can define a Dehn twist $\tau_{L}$ locally as $\tau_{L}=(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}\operatorname{Flow}^{1}_{X_{b}}$ where $b=b(\left\lVert p\right\rVert)$ is a function which equals 00 near $\left\lVert p\right\rVert=0$ and $\pi$ near $\left\lVert p\right\rVert=1$, $X_{b}$ is its Hamiltonian vector field $db=dp\wedge dq(\ast,X_{b})$, and $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ is the symplectomorphism induced by the antipodal map $q\mapsto-q$ on $\mathbb{S}^{n}$. In words, $\operatorname{Flow}^{1}_{X_{b}}$ applies the time $b(\left\lVert p\right\rVert)$ geodesic flow to (co)vectors of length $\left\lVert p\right\rVert$. In particular, when $\left\lVert p\right\rVert=1$, the time $1$-flow will send a vector over $q\in\mathbb{S}^{n}$ to a vector over $-q$ and the $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ correction ensures that $\tau_{L}$ is the identity near $\partial\mathbb{D}^{\ast}\mathbb{S}^{n}$. Extend $\tau_{L}$ to $F\setminus N_{F}(L)$ as the identity. The dependence on $b$ is irrelevant when we view $\tau_{L}\in\pi_{0}\operatorname{Symp}^{c}$. It can be seen by drawing pictures that when $\dim F=2$, $\tau_{L}$ coincides with the usual notion of a Dehn twist along a simple closed curve on a Riemann surface. The action of $\tau_{L}$ on $H_{\ast}(F)$ is given by the Picard-Lefschetz formula
(2.1.1) $(\tau_{L})h=\begin{cases}h&h\in H_{\neq n}(F),\\ h+(-1)^{\frac{1}{2}(n+1)(n+2)}(h\bullet[L])[L]&h\in H_{n}(F).\end{cases}$
In the cases $n=1,3$, $\tau_{L}^{2}$ is smoothly – but not symplectically – isotopic to the identity [Av21, Se99]. See [KRW23] for more on powers of Dehn twists.
A positive factorization of $\phi\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{F})$ is a tuple $\vec{L}=(L_{1},\cdots L_{m})$ of framed, exact Lagrangian spheres for which
@@
We say that $m$ is the length of the positive factoriation, written
@@
Two positive factorizations $\vec{L}$ and $\vec{L}^{\prime}=(L_{1}^{\prime},\cdots L_{m^{\prime}}^{\prime})$ of a $\phi$ are deformation equivalent if $m=m^{\prime}$ and for each $i$, $L_{i}$ is Hamiltonian isotopic to $L_{i}^{\prime}$, whence we write $\vec{L}=\vec{L}^{\prime}$. We are interested in such deformation equivalence classes of $\vec{L}$ as the provide instructions for building Weinstein manifolds of $\dim=\dim F+2$ as described in §2.3.
@@
A positive relation on $F$ is a pair of positive factorizations $\vec{L},\vec{L}^{\prime}$ of some $\phi$.
@@ 2.2. Known relations and counting twists
There are a few well-known ways to modify a positive factorization – and so determine positive relations:
@@
    The disjoint union relation states that $L\cap L^{\prime}=\emptyset\implies\tau_{L}\tau_{L^{\prime}}=\tau_{L^{\prime}}\tau_{L}$.
@@
    The conjugacy relation states that for any two $L$ and $L^{\prime}$, $\tau_{L}\tau_{L^{\prime}}=\tau_{\tau_{L}L^{\prime}}\tau_{L}$.
@@
    The braid relation states that if $L$ and $L^{\prime}$ intersect transversely in a single point, $\tau_{L}\tau_{L^{\prime}}\tau_{L}=\tau_{L^{\prime}}\tau_{L}\tau_{L^{\prime}}$.
@@
We seek to single out the positive factorizations which are not derived from braid groups and so are “not obvious”. Let $T=T(F,\beta_{F})$ be the free group on Hamiltonian isotopy classes of exact framed Lagrangian spheres. Consider the normal subgroup $B=B(F,\beta_{F})\subset T$ generated by $[L][L^{\prime}][L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ which are disjoint, $[L][L^{\prime}][L]^{-1}[\tau_{L}L^{\prime}]^{-1}$ for any $L,L^{\prime}$, and $[L][L^{\prime}][L][L^{\prime}]^{-1}[L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ intersecting transversely in $F$ a single point. We have canonical homomorphisms
$\tau:T\rightarrow\pi_{0}\operatorname{Symp}(F,\omega_{F}),\quad\tau[L]=\tau_{L},\quad p:T\rightarrow\mathbb{Z},\quad p[L]=1$
which factor through $T/B$. In words, $\tau$ sends formal products of exact Lagrangian spheres (and their inverses) to their products of their positive (and negative) Dehn twists and $p$ measures the sum of the exponents. So for a positive factorization $\vec{L}$, $p\vec{V}=|\vec{L}|$. By construction $\ker\tau$ gives all relations in the subgroup $\operatorname{im}\tau$ of compactly supported symplectomorphisms generated by Dehn twists.
Each positive factorization $\vec{L}$ of some $\phi$ determines elements of $T$ and $T/B$ by $[\vec{L}]=[L_{k}]\cdots[L_{1}]$. A positive relation $\vec{L},\vec{L}^{\prime}$ is non-braid-like if $[\vec{L}]\neq[\vec{L}^{\prime}]\in T/B$. It’s clear that if $|\vec{L}|\neq|\vec{L}^{\prime}|$, then the relation is non-braid-like. So examples of non-braid-like relations include the lantern relations and chain relations on punctured Riemann surfaces, cf. [OS04b, §15].
We are interested in finding non-braid like positive relations in the case $\dim F>2$. In a way which is made precise in [K22], the existence of such relations for $\dim F>2$ is more restrictive than the $\dim F=2$ case. Here is an easy way to see that relations in the $\dim F=2$ case (such as the lantern, which has $p=\pm 1$) cannot be straightforwardly generalized to higher dimensions.
@@
Suppose that $4$ divides $\dim F$ and $\vec{L}\in T$ is such that $\tau\vec{L}$ acts trivially on $H_{n}(F)$ (which of course must be satisfied if $\tau\vec{L}=\operatorname{Id}_{F}$). Then $p\vec{L}\in 2\mathbb{Z}$.
@@
Consider the action of a single $\tau_{L}$ on the determine line $\bigwedge^{top}H_{n}(F)$. Since $[L]$ is primitive, we can extend $[L]$ to a basis of $H_{n}(F)$ so that $\tau_{L}\in\operatorname{Aut}(H_{n})$ can be expressed as a matrix. The first row has $-1$ in the first entry (combining $[L]\bullet[L]=-\chi(\mathbb{S}^{n})=-2$ with Equation (2.1.1)) with the rest of the entries being zeros. In the $j$th row for $j>1$, Equation (2.1.1) determines the first entry, the $j$th entry is $1$, and all other entries are zero. So $\tau_{L}$ acts on the determinant line by $-1$. For a general $\vec{L}\in T$, $\tau\vec{L}$ acts on the determinant line by $(-1)^{p\vec{L}}$. ∎

### v2

@@ Abstract.
We describe symplectic mapping class relations between products of positive Dehn twists along Lagrangian spheres in Weinstein $4$-manifolds, all of which are affine $\mathbb{C}$ varieties. The relations are obtained by applying classification results for Fano $3$-folds and polarized $K3$ surfaces of small genus to a general methodology – finding pencil pairs.
@@ 1. Introduction
This article describes a general method for finding symplectic mapping class relations between products of positive Dehn twists on (affine algebraic) Weinstein manifolds $F$ of $\dim=2n$ by studying complex projective varieties of $\dim_{\mathbb{C}}=n+1$. The general method consists of finding pencil pairs and is known to experts. It recovers known mapping class relations on punctured Riemann surfaces such as the lantern relation [AS08, D87, J79], described in §3.3 and depicted in Figure 1. In formally defining pencil pairs and working out some basic techniques for finding them (see §4), we hope to inspire specialists in symplectic topology and algebraic geometry to further study the case $\dim F\geq 4$, for which few examples of such symplectic mapping class relations are known. Therefore our exposition will cover some relevant basics from both subject areas.
Our search for mapping class relations on high-dimensional Weinstein manifolds is motivated by recent advances in high-dimensional contact topology [Br23, BHH23, HH19]. In §2.3 we describe how these advances imply that relations between products of positive Dehn twists on Weinstein manifolds of any even dimension exist in abundance.
@@
Figure 1. The lantern relation states that the product $\tau_{c}\tau_{b}\tau_{a}$ of Dehn twists along circles on the left is boundary-relative isotopic to a product of Dehn twists along the circles on the right. Throughout, products represent compositions of maps, so $\tau_{c}\tau_{b}\tau_{a}$ means that we apply $\tau_{a}$, then $\tau_{b}$, and finally $\tau_{c}$. The surfaces depicted are oriented by a normal vector which sticks out of the page along their front halves.
In [O21] Oba finds a symplectic mapping class relation on a Weinstein $4$-manifold by applying the classification of del Pezzo surfaces [F80] to find a pencil pair. Our main application applies classification results for polarized $K3$ varieties [M87, SD74] and Fano $3$-folds [I77, I77, MM82] to find new pencil pairs, and so new examples of mapping class relations on Weinstein $4$-manifolds. Here is one such family of examples.
@@
Let $Z\subset\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{2}$ be the $K3$ surface cut out by a smooth complete intersection of $\deg=(1,1,2)$ and $\deg=(1,1,1)$ divisors. Let $(F,\beta_{F})$ be the Weinstein manifold given by the complement of a $\deg=(1,1,1)$ divisor in $Z$, with the symplectic structure determined by a $\deg=(1,1,1)$ line bundle. Then there is a compactly supported symplectomorphism $\tau_{\partial F}\in\pi_{0}\operatorname{Symp}^{c}(F,d\beta_{F})$ which can be factored into a product of $m$ Dehn twists along Lagrangian spheres in $(F,\beta_{F})$, where
$m=56,60,62,64,78.$
We found six such families of mapping class relations in total by applying the Fano $3$-fold and $K3$ surface classifications. One recovers the relation of [O21] (see §5.3). Because the number of Dehn twists in our relations are all distinct, they cannot be derived from well-known braid relations [KS02]. See §2.2. The $\tau_{\partial F}$ above is a fibered Dehn twist [CDvK14] and all mapping class relations coming from pencil pairs are determined by factorizations of such $\tau_{\partial F}$. See §2.4.
Every pencil pair gives rise to an infinite family of pencil pairs by a generally-applicable “cabling” result described in Theorem 4.2.2. Consequently, each collection of mapping class relations we have described above determines infinitely many mapping class relations. The proof of Theorem 4.2.2 relies on a general result concerning the symplectic topology of high-degree hypersurfaces in projective varieties, Theorem 4.2.1, which may be of independent interest.
@@ 1.1. Conventions
All algebraic varieties in this article are smooth and defined over $\mathbb{C}$. Throughout $\dim$ and $\dim_{\mathbb{C}}$ refer to real and complex dimension, respectively. On a closed manifold $X$ we use $\operatorname{PD}_{X}$ for Poincaré duality in either direction $H_{k}(X)\leftrightarrow H^{\dim X-k}(X)$. Cotangent disk bundles (determined by some Riemannian metric) over a smooth manifold $X$ will be denoted $\mathbb{D}^{\ast}X\subset T^{\ast}X$. On $T^{\ast}X$, the canonical $1$-form is denoted $pdq$ whose differential, denoted $dp\wedge dq$ is the canonical symplectic form. All other symplectic manifolds in this article will be complex projective varieties $X,Z,B,\mathcal{M}\subset\mathbb{CP}^{n}$ or affine varieties $F,W\subset\mathbb{C}^{n}$, with their symplectic structures inherited from the ambient space. For a function $h$ on a symplectic manifold with symplectic form $\omega$, the Hamiltonian vector field $X_{h}$ is defined by $dh=\omega(\ast,X_{h})$.
1.2. Acknowledgments
We thank Eduardo Alves da Silva, Yankı Lekili, and Takahiro Oba for interesting discussions. We’re especially grateful to Marco Golla for discussions around complex surfaces, reviewing an early draft of the article, and for helping to expand the list of examples in 3.3. We also thank our anonymous referees for their thoughtful feedback which has greatly improved the quality of this article. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034255. We thank Cofund MathInGreaterParis and the Fondation Mathématique Jacques Hadamard for supporting us as a member of the Laboratoire de Mathématiques d’Orsay at Université Paris-Saclay during the completion of the first draft of the article. Finally, we thank the CNRS for supporting us as a member of the Institut de Mathématiques de Jussieu-Paris Rive Gauche at Sorbonne Université.
@@ 2. Background
Here we cover background material and provide motivation for our interest in finding symplectic mapping class relations between products of positive Dehn twists.
@@ 2.1. Symplectic mapping class relations
An exact Lagrangian sphere $L\subset F$, where $\dim F=2n$, is a sphere $\mathbb{S}^{n}$ for which $\beta_{F}|_{TL}=df$ for some $f\in\mathcal{C}^{\infty}(L)$. A framing on such a sphere is an identification of the sphere with the boundary of a disk, up to homotopy. We give a quick review of Dehn twists about such $L$, which appear in nature as monodromies of Milnor fibers of the singularities $\sum_{j=0}^{n}z_{j}^{2}$ on $\mathbb{C}^{n+1}$ [Arn95]. According to the Weinstein neighborhood theorem, every exact Lagrangian sphere $L$ has a neighborhood of the form
$(N_{F}(L),\beta_{F})=(\mathbb{D}^{\ast}\mathbb{S}^{n},pdq+df),\quad f\in\mathcal{C}^{\infty}(\mathbb{D}^{\ast}\mathbb{S}^{n}).$
Assume the disk bundle is defined using the round metric of radius $1$. We define a Dehn twist $\tau_{L}$ locally as
$\tau_{L}=(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}\operatorname{Flow}^{1}_{X_{b}}$
where $b=b(\left\lVert p\right\rVert)$ is a function which equals 00 near $\left\lVert p\right\rVert=0$ and $\pi$ near $\left\lVert p\right\rVert=1$, $X_{b}$ is the Hamiltonian vector field, and $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ is the symplectomorphism induced by the antipodal map $q\mapsto-q$ on $\mathbb{S}^{n}$. In words, $\operatorname{Flow}^{1}_{X_{b}}$ applies the time $b(\left\lVert p\right\rVert)$ geodesic flow to (co)vectors of length $\left\lVert p\right\rVert$. In particular, when $\left\lVert p\right\rVert=1$, the time $1$-flow will send a vector over $q\in\mathbb{S}^{n}$ to a vector over $-q$ and the $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ correction ensures that $\tau_{L}$ is the identity near $\partial\mathbb{D}^{\ast}\mathbb{S}^{n}$. Extend $\tau_{L}$ to $F\setminus N_{F}(L)$ as the identity. The dependence on $b$ is irrelevant when we view $\tau_{L}\in\pi_{0}\operatorname{Symp}^{c}$. It can be seen by drawing pictures that when $\dim F=2$, $\tau_{L}$ coincides with the usual notion of a (right-handed) Dehn twist along a simple closed curve on a Riemann surface. The action of $\tau_{L}$ on $H_{\ast}(F)$ is given by the Picard-Lefschetz formula
(2.1.1) $(\tau_{L})h=\begin{cases}h+(-1)^{\frac{1}{2}(n+1)(n+2)}(h\bullet[L])[L]&h\in H_{n}(F),\\ h&\text{otherwise.}\\ \end{cases}$
Here $\bullet$ denotes the intersection pairing on $H_{n}(F)$. As a sanity check, we get
(2.1.2) $[L]\bullet[L]=-\chi(L),\quad(\tau_{L})[L]=(-1)^{n+1}[L]$
which is expected since $\tau_{L}$ acts as the antipodal map on $L\simeq\mathbb{S}^{n}$. In the cases $n=2,6$, $\tau_{L}^{2}$ is smoothly, but not symplectically, isotopic to the identity [Av21, Se99]. See [KRW23] for more on powers of Dehn twists.
A positive factorization of $\phi\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{F})$, denoted as $\vec{L}=(L_{1},\cdots L_{m})$, is a tuple of framed, exact Lagrangian spheres for which
@@
We say that $m$ is the length of the positive factorization, and write
@@
Two positive factorizations $\vec{L}$ and $\vec{L}^{\prime}=(L_{1}^{\prime},\cdots L_{m^{\prime}}^{\prime})$ of a $\phi$ are deformation equivalent if $m=m^{\prime}$ and for each $i$, $L_{i}$ is Hamiltonian isotopic to $L_{i}^{\prime}$, whence we write $\vec{L}=\vec{L}^{\prime}$. We are interested in such classes of $\vec{L}$ as they provide instructions for building Weinstein manifolds of $\dim=\dim F+2$ as described in §2.3.
@@
A positive relation on $F$ is a pair of positive factorizations $\vec{L},\vec{L}^{\prime}$ of some $\phi$, with $\vec{L}\neq\vec{L}^{\prime}$.
@@ 2.2. Known relations and counting twists
There are a few well-known ways to modify a positive factorization, and so determine positive relations.
@@
    The disjoint union relation: If $L\cap L^{\prime}=\emptyset$, then $\tau_{L}\tau_{L^{\prime}}=\tau_{L^{\prime}}\tau_{L}$.
@@
    The conjugacy relation: For any two $L$ and $L^{\prime}$, $\tau_{L}\tau_{L^{\prime}}=(\tau_{\tau_{L}L^{\prime}})\tau_{L}$.
@@
    The braid relation: If $L$ and $L^{\prime}$ intersect transversely in a single point, $\tau_{L}\tau_{L^{\prime}}\tau_{L}=\tau_{L^{\prime}}\tau_{L}\tau_{L^{\prime}}$.
@@
We seek to single out the positive factorizations which are not derived from braid groups and so are “not obvious”. Let $T=T(F,\beta_{F})$ be the free group on Hamiltonian isotopy classes of exact framed Lagrangian spheres. Consider the normal subgroup $B=B(F,\beta_{F})\subset T$ generated by
  1. (1)
    $[L][L^{\prime}][L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ which are disjoint,
  2. (2)
    $[L][L^{\prime}][L]^{-1}[\tau_{L}L^{\prime}]^{-1}$ for any $L,L^{\prime}$,
  3. (3)
    and $[L][L^{\prime}][L][L^{\prime}]^{-1}[L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ intersecting transversely in $F$ a single point.
We have canonical homomorphisms
(2.2.1) $\begin{gathered}\tau:T\rightarrow\pi_{0}\operatorname{Symp}(F,\omega_{F}),\quad\tau[L]=\tau_{L},\\ p:T\rightarrow\mathbb{Z},\quad p[L]=1\end{gathered}$
which factor through $T/B$. In words, $\tau$ sends formal products of exact Lagrangian spheres (and their inverses) to products of their positive (and negative) Dehn twists while $p$ measures the sum of the exponents. Therefore for a positive factorization $\vec{L}$, $p[\vec{L}]=|\vec{L}|$. By construction, $\ker\tau$ gives all relations in the subgroup $\operatorname{im}\tau$ of compactly supported symplectomorphisms generated by Dehn twists.
Each positive factorization $\vec{L}$ of some compactly supported symplectomorphism $\phi$ yields
$[\vec{L}]=[L_{k}]\cdots[L_{1}]$
which can be viewed as an element of either $T$ or $T/B$. A positive relation $\vec{L},\vec{L}^{\prime}$ is non-braid-like if $[\vec{L}]\neq[\vec{L}^{\prime}]\in T/B$. It’s clear that if $|\vec{L}|\neq|\vec{L}^{\prime}|$, then the relation is non-braid-like. Examples of non-braid-like relations include the lantern and chain relations on punctured Riemann surfaces, cf. [OS04b, §15].
We are interested in finding non-braid-like positive relations when $\dim F>2$. In a way which is made precise in [K22], the existence of such relations is more restrictive than the $\dim F=2$ case. The following lemma gives an easy way to see that relations in the $\dim F=2$ case (such as the lantern, which has $p=\pm 1$) cannot be straightforwardly generalized to higher dimensions.
@@
Following the notation of Equation (2.2.1), suppose that $4$ divides $\dim F=2n$ and $\vec{L}\in T$ is such that $\tau\vec{L}$ acts trivially on $H_{n}(F)$ (which of course must be satisfied if $\tau\vec{L}=\operatorname{Id}_{F}$). Then $p\vec{L}\in 2\mathbb{Z}$.
@@
Consider the action of a single Dehn twist $\tau_{L}$ on the determinant line $\ell=\bigwedge^{top}H_{n}(F)$. Since $n$ is even, $[L]\bullet[L]=-2$ which isn’t divisible by a squared integer, implying that $[L]\in H_{n}(F)$ is primitive. Hence we can extend $[L]$ to a basis of $H_{n}(F)$ with respect to which $\tau_{L}\in\operatorname{Aut}(H_{n})$ can be expressed as a matrix. The first row has $-1$ in the first entry by Equation (2.1.2), with the rest of the entries being zeros. In the $j$th row for $j>1$, Equation (2.1.1) determines the first entry, the $j$th entry is $1$, and all other entries are zero. Therefore $\tau_{L}$ acts on $\ell$ by $-1$. For a general $\vec{L}\in T$, $\tau\vec{L}$ acts on the determinant line by $(-1)^{p\vec{L}}$. Since we’re assuming that $\tau\vec{L}$ acts trivially on $H_{n}(F)$, it must act trivially on $\ell$ and $2$ must divide $p\vec{L}$. ∎